To whom it may concern,

Hello, I need help with this question and seeing if I am doing the work and drawings correctly with units.

Question:

"Two electrons are 2.40 m. apart. Another electron is shot from infinity and comes to rest midway between the two. What must its initial velocity be (in meters/second)? "

Let me know when you can. Thank you!
Answer #1

conserving energy at intial postion then final.

If any query please comment if you like my solution give me thumbs up
Answer #2

Given:

  • Two fixed electrons are 2.40 meters apart.

  • A third electron is shot from infinity and comes to rest midway between them (i.e., at 1.20 m from each fixed electron).

  • Charge of an electron, e=1.602×1019C.

  • Mass of an electron, me=9.109×1031kg.

  • Coulomb's constant, k=8.988×109Nm2C2.


Step 1: Calculate the Electric Potential Energy (Final State)

When the third electron comes to rest at the midpoint, its total potential energy is due to repulsion from the two fixed electrons:

Ufinal=2×(ke2r)

where r=1.20m.

Ufinal=2×((8.988×109)(1.602×1019)21.20)Ufinal=2×(1.921×1028J)=3.842×1028J


Step 2: Electric Potential Energy at Infinity (Initial State)

At infinity (r), the potential energy is:

Uinitial=0J


Step 3: Apply Energy Conservation

The electron's initial kinetic energy is converted to potential energy at the midpoint:

Kinitial=UfinalUinitial12mev2=3.842×1028J

Solve for velocity v:

v=2×3.842×10289.109×1031v=8.436×102=29.04m/s


Final Answer:

The initial velocity of the electron must be 29.04m/s.


answered by: anonymous